OpenStudy (anonymous):

x+y=2. solve for x and y

OpenStudy (alexwee123):

x+y=2 subtract either x or y on both side depending on what you want and in this case you need both! so... y=2-x and x=2-y

OpenStudy (alexwee123):

questions? :o

OpenStudy (anonymous):

since you have one equation and two unknowns, then there is no 1 solution for the equation. all you can do is specify x in terms of y or y in terms of x e.g. x = 2-y or y = 2-x